If a certain number x is subtracted from 8, it will be equal to two less than four times the same number. The correct equation for the above sentence is: [1] x − 8 = 4x − 2 [2] 8 − x = 4 − 2x [3] 8 − x = 4x − 2 [4] x − 8 = 2x − 4


1
Expert's answer
2022-02-22T08:46:30-0500

If a certain number x is subtracted from 8, it will be equal to two less than four times the same number, or:

8 - x = 2 - 4x

or

x - 8 = 4x - 2.

So, the correct answer is [1] x − 8 = 4x − 2.